# Permutation-Invariant and Orientation-Aware Dataset Distillation for 3D Point Clouds

### v1

@@ Abstract
We should collect large amount of data to train deep neural networks for various applications. Recently, the dataset distillation for images and texts has been attracting a lot of attention, that reduces the original dataset to a synthetic dataset while preserving essential task-relevant information. However, 3D point clouds distillation is almost unexplored due to the challenges of unordered structures of points. In this paper, we propose a novel distribution matching-based dataset distillation method for 3D point clouds that jointly optimizes the geometric structures of synthetic dataset as well as the orientations of synthetic models. To ensure the consistent feature alignment between different 3D point cloud models, we devise a permutation invariant distribution matching loss with the sorted feature vectors. We also employ learnable rotation angles to transform each syntheic model according to the optimal orientation best representing the original feature distribution. Extensive experimental results on widely used four benchmark datasets, including ModelNet10, ModelNet40, ShapeNet, and ScanObjectNN, demonstrate that the proposed method consistently outperforms the existing methods.
@@ 1 Introduction
With the increasing demand for large-scale datasets, dataset distillation [21] has gained significant attention across various applications of deep learning. The goal of dataset distillation is to reduce the size of an original dataset into a synthetic dataset that preserves the essential task-relevant information of the original dataset. Therefore, the models trained on the reduced synthetic dataset are encouraged to achieve comparable performance to those trained on the original dataset. Existing dataset distillation methods [3, 6, 8, 11, 29, 28, 30, 17, 26, 5, 9] can be broadly classified into gradient matching, trajectory matching, and distribution matching approaches. The gradient matching and trajectory matching methods aim to ensure that the synthetic dataset produces similar optimization dynamics to the original dataset. To this end, the gradient matching method [29] minimizes the difference between the gradients computed on the synthetic and original datasets. However, these methods potentially overlook long-term dependencies in the training process. Rather than comparing individual gradients, the trajectory matching methods [3, 6, 8, 11] encourage the models trained on a synthetic dataset follow similar optimization trajectories to those trained on the original dataset. However, these methods train the networks while optimizing the synthetic dataset significantly increasing the computational complexity. To alleviate the computational burden in dataset distillation, the distribution matching techniques [28, 30, 17, 26, 5] have been introduced. They randomly initialize the networks without training to extract the features, and compare the feature distributions between the original and synthetic datasets.
Figure 1: The concept of the proposed dataset distillation method for 3D point clouds. (a) Permutation-invariant feature mathcing. (b) Orientation optimization.
While the dataset distillation has been extensively studied for structured data, such as images [29, 3, 28] and texts [10, 12, 13], that for 3D point clouds is almost unexplored. Recently, PCC [27] applied an existing gradient matching method of DC [29] to tackle the data distillation problem of 3D point clouds, but it did not consider the inherent characteristics of 3D point clouds and still suffers from the high computational cost. Unlike structured data, 3D point clouds consist of unordered and irregularly distributed points in 3D space. Specifically, inconsistent indexing of points makes it challenging to develop efficient dataset distillation techniques. Furthermore, 3D models often exhibit different orientation from one another, that also degrades the distillation performance.
@@
In this paper, we propose a joint optimization framework of Permutation Invariant Distribution Matching (PIDM) and orientation optimization for dataset disllation of 3D point clouds. A major challenge in this setting is achieving effective feature alignment despite the unordered nature of point clouds and their arbitrary orientations. As illustrated in Figure 1(a), the PIDM loss addresses the issue of unordered structure by sorting the elements within each feature vector according to their sizes before computing the distance between two feature distributions. To further enhance the feature alignment, as shown in Figure 1(b), We also employ learnable parameters of the rotation angles to estimate optimal orientation of 3D models, while optimizing the synthetic dataset. Experimental results demonstrate that the proposed method optimizes both the synthetic dataset and the orientation of 3D models simultaneously outperforming the existing methodologies.
@@
  • To the best of our knowledge, we are the first to proposed a distribution matching-based dataset distillation method for 3D point clouds, that jointly optimizes the synthetic dataset and the orientation of 3D models simultaneously.
  • We devised PIDM loss to ensure permutation invariant feature alignment for 3D point clouds with inconsistent orders of points.
  • We validated proposed method through comprehensive experiments on four benchmark datasets of 3D point clouds classification, and showed the superiority of the proposed method over the existing dataset distillation techniques.
@@ 2 Related Works
Figure 2: The overall framework of the proposed 3D point clouds dataset distillation method.
@@ 3 Methodology
We propose a novel dataset distillation method for 3D point clouds that first addressing the key challenges of unordered structure and rotational variation of 3D point clouds. We perform PIDM that effectively aligns the features of unordered points across different 3D models. To handle the rotational variation, we also estimate optimal orientation while generating the synthetic dataset. Figure 2 shows the overall framework of the proposed method.
3.1 Problem Formulation
The objective of dataset distillation is to compress the task-relavant information in the original dataset ${\boldsymbol{\mathcal{T}}}=\left\{{{\bf t}}_{i}\right\}$ and generate a much smaller synthetic dataset ${\boldsymbol{\mathcal{S}}}=\left\{{{\bf s}}_{i}\right\}$, $|{\boldsymbol{\mathcal{S}}}|\ll|{\boldsymbol{\mathcal{T}}}|$, where the models trained on ${\boldsymbol{\mathcal{S}}}$ achieve close performance to that trained on ${\boldsymbol{\mathcal{T}}}$. Given a 3D point cloud sample ${{\bf p}}$ following a real data distribution with the corresponding class label $l$, the optimal synthetic dataset ${\boldsymbol{\mathcal{S}}}^{\star}$ can be obtained via
@@ 3.1 Problem Formulation
where ${\mathcal{L}}$ denotes a task-specific loss function, such as the cross-entropy loss, and $\phi_{{\boldsymbol{\mathcal{T}}}}$ and $\phi_{{\boldsymbol{\mathcal{S}}}}$ represent the models trained on ${\boldsymbol{\mathcal{T}}}$ and ${\boldsymbol{\mathcal{S}}}$, respectively.
3.2 Permutation Invariant Distribution Matching
We basically perform the distribution matching by minimizing the discrepancy between the feature distributions derived from the original and synthetic datasets, respectively, such that
@@ 3.2 Permutation Invariant Distribution Matching
where $D$ is a distance function, and $\phi$ denotes a randomly initialized untrained network. Note that we simply use a randomly initialized network without training and effectively alleviate the computational overhead in dataset distillation. The Maximum Mean Discrepancy (MMD) [7] loss $\mathcal{L}_{\text{MMD}}$ is often used to measure $D$, defined as
@@
where ${K}(\cdot,\cdot)$ is the kernel function, where we used the Gaussian kernel in this work, given by
${K}({{\boldsymbol{\mathcal{T}}},\boldsymbol{\mathcal{S}}})=\frac{1}{|\boldsymbol{\mathcal{T}}|\cdot|\boldsymbol{\mathcal{S}}|}\sum_{{{{\bf t}}\in\boldsymbol{\mathcal{T}}}}\sum_{{{\bf s}}\in\boldsymbol{\mathcal{S}}}\text{exp}\left(-\frac{||\phi({{\bf t}})-\phi({{\bf s}})||^{2}}{2\sigma}\right).$ (4)
Note that the spatial relationships among the pixels are well structured and consistent across different images, but the points in 3D space are usually indexed in different orders across multiple 3D models. Therefore, the conventional distribution matching methods cannot be directly applied to 3D point clouds, since they assume aligned features between two compared models. To overcome this issue, we propose PIDM that performs the feature sorting in advance when comparing the feature distributions between the original and synthetic datasets. Specifically, given a 3D point cloud model ${{\bf p}}$, we extract point-wise features by using $\phi$ as
@@
where $N$ denotes the number of points in ${{\bf p}}$, and $C$ is the number of feature channels. We first rearrange the elements of ${\bf f}_{i}$ in descending order of their sizes to have the sorted feature vector $\tilde{{\bf f}}_{i}$. Then we have the sorted features as
@@
Note that ${{\bf t}}$ in the original dataset $\boldsymbol{\mathcal{T}}$ and ${{\bf s}}$ in the synthetic dataset $\boldsymbol{\mathcal{S}}$ exhibit different orders of points in general, and therefore the features in ${\phi}({{\bf t}})$ and ${\phi}({{\bf s}})$ are inconsistenly aligned with each other disrupting the desired behavior of comparison in (4). On the contrary, $\tilde{\phi}({{\bf t}})$ and $\tilde{\phi}({{\bf s}})$ yield well aligned features invariant to the permutation variation, and therefore, we compute the refined Gaussian kernel with the sorted features, given by
@@
$\tilde{K}({{\boldsymbol{\mathcal{T}}},\boldsymbol{\mathcal{S}}})=\frac{1}{|\boldsymbol{\mathcal{T}}|\cdot|\boldsymbol{\mathcal{S}}|}\sum_{{{{\bf t}}\in\boldsymbol{\mathcal{T}}}}\sum_{{{\bf s}}\in\boldsymbol{\mathcal{S}}}\text{exp}\left(-\frac{||\tilde{\phi}({{\bf t}})-\tilde{\phi}({{\bf s}})||^{2}}{2\sigma}\right).$ (7)
@@
Accordingly, we devise the PIDM loss $\mathcal{L}_{\text{PIDM}}$ from $\mathcal{L}_{\text{MMD}}$, given by
$\mathcal{L}_{\text{PIDM}}({\boldsymbol{\mathcal{T}},\boldsymbol{\mathcal{S}}})={K}({\boldsymbol{\mathcal{T}},\boldsymbol{\mathcal{T}}})+{K}({\boldsymbol{\mathcal{S}},\boldsymbol{\mathcal{S}}})-2\tilde{K}({\boldsymbol{\mathcal{T}},\boldsymbol{\mathcal{S}}}).$ (8)
By using $\mathcal{L}_{\text{PIDM}}$ instead of $\mathcal{L}_{\text{MMD}}$, we improve the performance of datasets distillation for 3D point clouds regardless of the orders of points.

### v2

# Dataset Distillation of 3D Point Clouds via Distribution Matching

@@ Abstract
Large-scale datasets are usually required to train deep neural networks, but it increases the computational complexity hindering the practical applications. Recently, dataset distillation for images and texts has been attracting a lot of attention, that reduces the original dataset to a synthetic dataset to alleviate the computational burden of training while preserving essential task-relevant information. However, the dataset distillation for 3D point clouds remains largely unexplored, as the point clouds exhibit fundamentally different characteristics from that of images, making the dataset distillation more challenging. In this paper, we propose a distribution matching-based distillation framework for 3D point clouds that jointly optimizes the geometric structures as well as the orientations of the synthetic 3D objects. To address the semantic misalignment caused by unordered indexing of points, we introduce a Semantically Aligned Distribution Matching loss computed on the sorted features in each channel. Moreover, to address the rotation variation, we jointly learn the optimal rotation angles while updating the synthetic dataset to better align with the original feature distribution. Extensive experiments on widely used benchmark datasets demonstrate that the proposed method consistently outperforms existing dataset distillation methods, achieving superior accuracy and strong cross-architecture generalization.
@@ 1 Introduction
With the increasing demand for large-scale training datasets, the cost of model retraining from scratch has become a major challenge leading to the research of dataset distillation [21]. The objective of dataset distillation is to produce a significantly smaller synthetic dataset from a large original dataset, that preserves the essential task-relevant information contained in the original dataset. Hence the models trained on the reduced synthetic dataset are encouraged to achieve comparable performance to those trained on the original dataset. Existing dataset distillation methods [3, 6, 8, 11, 29, 28, 30, 17, 26, 5, 9] can be broadly classified into gradient matching, trajectory matching, and distribution matching approaches. The gradient matching and trajectory matching methods aim to ensure that the synthetic dataset produces similar optimization dynamics to the original dataset. To this end, the gradient matching method [29] minimizes the difference between the gradients computed on the synthetic and original datasets. However, it potentially overlooks long-term dependencies in the training process. Rather than comparing individual gradients, the trajectory matching methods [3, 6, 8, 11] encourage the models trained on the synthetic dataset follow similar optimization trajectories to those trained on the original dataset. Note that these methods train the networks while optimizing the synthetic dataset, and significantly increase the computational complexity. To alleviate the computational burden in dataset distillation, the distribution matching techniques [28, 30, 17, 26, 5] have been introduced. They randomly initialize the networks without training to extract the features, and compare the feature distributions between the original and synthetic datasets.
While dataset distillation has been extensively studied for structured data such as images [29, 3, 28] and texts [10, 12, 13], its application to 3D point clouds remains almost unexplored. PCC [27] simply applied an existing distillation method for images [29] to the 3D point clouds without considering inherent characteristics of 3D point clouds, however it still suffers from high computational complexity. Unlike the structured data, 3D point clouds consist of unordered and irregularly distributed points in 3D space. Therefore, semantically similar regions across different 3D models are often associated with inconsistent orders (indices of points), that makes direct feature comparison incorrect worsening the performance of distribution matching based dataset distillation. We refer to this issue as semantic misalignment. Moreover, the datasets of 3D point clouds also suffers from rotational variation. 3D point clouds are often captured or synthesized under arbitrary poses due to the absence of canonical orientations. Such rotational difference causes the objects of the same class yield different features from one another, significantly increasing the intra-class variability. As a result, it becomes difficult to construct a representative synthetic dataset that faithfully follows the feature distribution of the original dataset.
@@
In this paper, we propose an optimization framework combining Semantically Aligned Distribution Matching (SADM) and orientation optimization for dataset distillation of 3D point clouds. A major challenge in this setting is achieving effective feature alignment despite the unordered nature of point clouds and their arbitrary orientations. The SADM loss addresses the issue of misaligned semantic structures by sorting the point-wise feature values within each channel before computing the distance between two feature distributions. In parallel, we employ learnable rotation parameters to address the orientation variation by estimating the optimal poses of 3D models while generating the synthetic dataset. By simultaneously enforcing semantic consistency and orientation alignment, the proposed method achieves superior performance of dataset distillation compared with the existing methods, as demonstrated by extensive experiments on standard 3D point cloud classification benchmarks.
@@
  • To the best of our knowledge, we are the first to propose a distribution matching-based dataset distillation method of 3D point clouds, that jointly optimizes the shapes and orientations of synthetic dataset.
  • We devised the SADM loss on the sorted features in each channel to preserve the semantic alignment between compared 3D objects.
  • We validated the proposed method through extensive experiments on the four benchmark datasets widely used for 3D point clouds classification, and showed the superiority of the proposed method over the existing dataset distillation techniques.
@@ 3 Methodology
Figure 1: The overall framework of the proposed dataset distillation method for 3D point clouds.
We propose a novel dataset distillation method for 3D point clouds that first addresses the key challenges of semantic misalignment and rotational variation. We perform SADM that effectively aligns the features of semantically consistent structures across different 3D models. To handle the rotational variation, we also estimate optimal orientations while generating the synthetic dataset. Figure 1 shows the overall framework of the proposed method.
3.1 Preliminaries
Problem Definition. The objective of dataset distillation is to compress the task-relavant information in the original dataset ${\boldsymbol{\mathcal{T}}}=\left\{{{\bf t}}_{i}\right\}_{i=1}^{|\mathcal{T}|}$ and generate a much smaller synthetic dataset ${\boldsymbol{\mathcal{S}}}=\left\{{{\bf s}}_{i}\right\}_{i=1}^{|\mathcal{S}|}$, where $|{\boldsymbol{\mathcal{S}}}|\ll|{\boldsymbol{\mathcal{T}}}|$, such that a model trained on ${\boldsymbol{\mathcal{S}}}$ achieves close performance to that trained on ${\boldsymbol{\mathcal{T}}}$. Given a 3D point cloud sample ${{\bf p}}$ following a real data distribution with the corresponding class label $l$, the optimal synthetic dataset ${\boldsymbol{\mathcal{S}}}^{\star}$ can be obtained via
@@ 3.1 Preliminaries
where ${\mathcal{L}}$ denotes a task-specific loss function, such as the cross-entropy loss, and $\phi_{{\boldsymbol{\mathcal{T}}}}$ and $\phi_{{\boldsymbol{\mathcal{S}}}}$ represent the models to estimate the class label which are trained on ${\boldsymbol{\mathcal{T}}}$ and ${\boldsymbol{\mathcal{S}}}$, respectively.
Distribution Matching. Distribution matching (DM) strategy focuses on aligning the feature distributions derived from the original and synthetic datasets, respectively, via
@@
where $\phi$ is the feature extractor and $D$ denotes a distance function. We employ a randomly initialized, untrained network, which has been demonstrated to sufficiently capture the structural information for distribution alignment [28]. Also, the Maximum Mean Discrepancy (MMD) [7] loss $\mathcal{L}_{\text{MMD}}$ is often used as $D$, given by
@@
where ${K}(\cdot,\cdot)$ is a kernel function. We used the Gaussian kernel in this work as
${K}({\boldsymbol{\mathcal{T}}},{\boldsymbol{\mathcal{S}}})=\frac{1}{|\boldsymbol{\mathcal{T}}|\cdot|\boldsymbol{\mathcal{S}}|}\sum_{{{\bf t}}\in\boldsymbol{\mathcal{T}}}\sum_{{{\bf s}}\in\boldsymbol{\mathcal{S}}}\exp\left(-\frac{\|\phi({{\bf t}})-\phi({{\bf s}})\|^{2}}{2\sigma}\right).$ (4)
3.2 Semantically Aligned Distribution Matching
Figure 2: Visualization of the points corresponding to the largest (top), 200th largest (middle), and 500th largest (bottom) features in each channel.
Whereas the pixels of image exhibit well structured spatial relationships with one another and consistently indexed across different images, the points in 3D point clouds are unordered with different indices across different models. Therefore, the conventional distribution matching methods cannot be directly applied to 3D point clouds, since the features of semantically similar structures are not aligned between two compared models. Motivated by the importance of preserving semantic correspondence, we investigate the relationship between the feature values and semantic significance in 3D point clouds. We first extracted point-wise features using a randomly initialized network $\phi$, and sorted the feature values for each of 1024 channels according to their size. Figure 2 visualizes the points corresponding to the largest (top), 200th largest (middle), and 500th largest (bottom) features, respectively. We observe that, even without model training, the points associated with the largest features consistently capture semantically meaningful regions, such as edges and corners, across different classes. In contrast, the points yielding low-ranked features are distributed around less informative regions. This suggests that the features reflect the relative importance of points in characterizing the structures of 3D objects, and the points with similar orders of sorted features tend to capture semantically related regions across different 3D models.
In order to preserve the semantic correspondence across different 3D models, we propose a semantically aligned distribution matching (SADM) loss. Specifically, given a 3D point cloud object ${{\bf p}}$, we extract the features using $\phi$ via
@@ 3.2 Semantically Aligned Distribution Matching
where $N$ denotes the number of points in ${{\bf p}}$, $C$ is the number of feature channels, and ${\bf f}_{i}=\left\{f_{i,1},f_{i,2},...,f_{i,N}\right\}$ represents the feature vector of the $i$-th channel where $f_{i,j}$ is the feature value of the $j$-th point. We then perform channel-wise sorting to the extracted features of $f_{i,j}$’s in the descending order of their sizes, and obtain the sorted feature vector $\tilde{{\bf f}}_{i}=\left\{\tilde{f}_{i,1},\tilde{f}_{i,2},...,\tilde{f}_{i,N}\right\}$ such that $\tilde{f}_{i,j}\geq\tilde{f}_{i,j+1}$. Then we have the set of the sorted feature vectors as
@@
Note that ${{\bf t}}$ in the original dataset $\boldsymbol{\mathcal{T}}$ and ${{\bf s}}$ in the synthetic dataset $\boldsymbol{\mathcal{S}}$ exhibit different orders of points in general, and therefore the features in ${\phi}({{\bf t}})$ and ${\phi}({{\bf s}})$ are inconsistently aligned with each other worsening the desired behavior of comparison in (3). On the contrary, the sorted features $\tilde{\phi}({{\bf t}})$ and $\tilde{\phi}({{\bf s}})$ exhibit consistent ordering of features in each channel reflecting their relative semantic importance, and thus facilitates reliable feature comparison across different 3D objects. To measure the discrepancy between the sorted features, we redefine the Gaussian kernel as
@@
$\tilde{K}({{\boldsymbol{\mathcal{T}}},\boldsymbol{\mathcal{S}}})=\frac{1}{|\boldsymbol{\mathcal{T}}|\cdot|\boldsymbol{\mathcal{S}}|}\sum_{{{{\bf t}}\in\boldsymbol{\mathcal{T}}}}\sum_{{{\bf s}}\in\boldsymbol{\mathcal{S}}}\exp\left(-\frac{||\tilde{\phi}({{\bf t}})-\tilde{\phi}({{\bf s}})||^{2}}{2\sigma}\right).$ (7)
@@
Then we devise the loss $\mathcal{L}_{\alpha}$ from $\mathcal{L}_{\text{MMD}}$ (3), given by
$\mathcal{L}_{\alpha}(\boldsymbol{\mathcal{T}},\boldsymbol{\mathcal{S}})=\tilde{K}(\boldsymbol{\mathcal{T}},\boldsymbol{\mathcal{T}})+\tilde{K}(\boldsymbol{\mathcal{S}},\boldsymbol{\mathcal{S}})-2\tilde{K}(\boldsymbol{\mathcal{T}},\boldsymbol{\mathcal{S}}).$ (8)
We additionally employ $\mathcal{L}_{\beta}$ to boost the role of the most significant feature in each channel, defined as
$\mathcal{L}_{\beta}(\boldsymbol{\mathcal{T}},\boldsymbol{\mathcal{S}})=\tilde{K}_{\text{top}}(\boldsymbol{\mathcal{T}},\boldsymbol{\mathcal{T}})+\tilde{K}_{\text{top}}(\boldsymbol{\mathcal{S}},\boldsymbol{\mathcal{S}})-2\tilde{K}_{\text{top}}(\boldsymbol{\mathcal{T}},\boldsymbol{\mathcal{S}}),$ (9)
where $\tilde{K}_{\text{top}}(\cdot,\cdot)$ denotes the Gaussian kernel computed with only the most largest feature in each channel. The SADM loss is then formulated as a weighted sum of the two losses, given by
$\mathcal{L}_{\text{SADM}}(\boldsymbol{\mathcal{T}},\boldsymbol{\mathcal{S}})=\lambda_{1}\mathcal{L}_{\alpha}(\boldsymbol{\mathcal{T}},\boldsymbol{\mathcal{S}})+\lambda_{2}\mathcal{L}_{\beta}(\boldsymbol{\mathcal{T}},\boldsymbol{\mathcal{S}}),$ (10)
where $\lambda_{1},\lambda_{2}$ are the weighting parameters. By using $\mathcal{L}_{\text{SADM}}$ instead of $\mathcal{L}_{\text{MMD}}$, we facilitate reliable matching of feature distributions considering sematic structures of 3D point clouds, thereby improving the performance dataset distillation.
3.3 Estimation of Optimal Rotations
3D objects usually exhibit different orientations from each other. Therefore, while generating optimal 3D objects in the synthetic dataset in terms of their geometric shapes, we also estimate their optimal rotations best representing various orientations of 3D objects in the original dataset. In practice, we introduce three rotation angles, $\theta_{x}$, $\theta_{y}$, and $\theta_{z}$, corresponding to rotations around the $x$-, $y$-, and $z$-axes, respectively. These angles are treated as learnable parameters, allowing the orientation of each synthetic 3D object in ${\boldsymbol{\mathcal{S}}}$ to be adaptively adjusted during dataset distillation. Therefore, instead of minimizing $\mathcal{L}_{\text{SADM}}({\boldsymbol{\mathcal{T}},\boldsymbol{\mathcal{S}}})$ in (10), we minimize $\mathcal{L}_{\text{SADM}}({\boldsymbol{\mathcal{T}},\hat{\boldsymbol{\mathcal{S}}}})$, where $\hat{\boldsymbol{\mathcal{S}}}$ is the synthetic dataset whose 3D objects are rotated toward the estimated optimal orientations.
Note that we optimize the synthetic 3D objects as well as their rotation parameters simultaneously, during the dataset distillation process. Specifically, at each iteration, we randomly initialize the network parameters and construct the synthetic dataset $\boldsymbol{\mathcal{S}}$ by randomly selecting samples from the original dataset. We then form $\hat{\boldsymbol{\mathcal{S}}}$ by rotating the objects in $\boldsymbol{\mathcal{S}}$ according to the angles $\theta_{x}$, $\theta_{y}$, and $\theta_{z}$, initially set to zero. For each class, a mini-batch is sampled from $\boldsymbol{\mathcal{T}}$, with a batch size of 8 per class. The corresponding synthetic mini-batch is sampled from $\hat{\boldsymbol{\mathcal{S}}}$, with the batch size determined by the number of point cloud objects per class (PPC). The joint optimization process iteratively updates both the geometric structure of the synthetic dataset and their rotation parameters by minimizing $\mathcal{L}_{\text{SADM}}({\boldsymbol{\mathcal{T}},\hat{\boldsymbol{\mathcal{S}}}})$. This ensures that the synthetic dataset preserves the geometric characteristics of the original dataset while aligning their orientations more effectively. The benefit of this joint optimization is justified in the following proposition.
Proposition 1. Jointly optimizing the synthetic dataset $\boldsymbol{\mathcal{S}}$ and the rotation parameters $\boldsymbol{\theta}=(\theta_{x},\theta_{y},\theta_{z})$ guarantees a lower or equal loss to that of optimizing $\boldsymbol{\mathcal{S}}$ alone.
$\min_{\boldsymbol{\mathcal{S}},\boldsymbol{\theta}}\mathcal{L}_{\text{SADM}}(\boldsymbol{\mathcal{T}},\mathcal{R}_{\boldsymbol{\theta}}(\boldsymbol{\mathcal{S}}))\leq\min_{\boldsymbol{\mathcal{S}}}\mathcal{L}_{\text{SADM}}(\boldsymbol{\mathcal{T}},\boldsymbol{\mathcal{S}}),$ (11)
where $\mathcal{R}_{\boldsymbol{\theta}}$ denotes the rotation operator according to $\boldsymbol{\theta}$.
Proof Sketch. The proof is provided in Appendix A. Allowing optimization over both $\boldsymbol{\mathcal{S}}$ and $\boldsymbol{\theta}$ enlarges the feasible set compared to optimizing only $\boldsymbol{\mathcal{S}}$. The feasible set for the joint optimization includes the feasible set of the original optimization as a subset, since optimizing only $\boldsymbol{\mathcal{S}}$ is equivalent to the special case of the joint optimization with the fixed $(\theta_{x},\theta_{y},\theta_{z})=(0,0,0)$. Consequently, the minimum loss achieved by jointly optimizing over $\boldsymbol{\mathcal{S}}$ and $\boldsymbol{\theta}$ must be less than or equal to the minimum loss achieved by optimizing only $\boldsymbol{\mathcal{S}}$. $\square$
Table 1: Comparison of quantitative performance. DC, DM, and the proposed method were initialized with random selection [16] for fair comparison. The best and the second best scores are highlighted in bold and underlined, respectively.